A 6-faced fair dice is rolled repeatedly till 1 appears for the first time. The probability that the dice is rolled for even number of time is________.

Grade:11

2 Answers

Aman Bansal
11 years ago

Dear Shiva,

Pr = (6P(r-1) * (r-1)) / 6^r
where the 6P(r-1) is the permutation function.
This gives
P1 = 0
P2 = 1/6
P3 = 10/36 = 5/18
P4 = 60/216 = 5/18
P5 = 5 / 27
P6 = 25/324
P7 = 5/324
P8 = 0
P9 = 0
P10 = 0

b) 1
The sum is the same as saying "What is the probability that the first time you have a repeated number is on the first ten rolls?" Well, we can only roll six times without getting a repeat, in which case the 7th roll would be our last. Therefore we ALWAYS have to get a repeated number in the first ten rolls, so the probability is 1.

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

upender surepally
11 years ago

5/11